On February 1, 2020, Oriole Contractors agreed to construct a building at a contract price of $3,130,000. The total estimated construction costs would be $1,555,000 and the project would be finished in 2022. Information relating to the costs and billings for this contract is as follows:

2020 2021 2022
Total costs incurred to date $ 544,250 $ 903,650 $ 2,380,000
Estimated costs to complete 1,010,750 801,350 -0-
Customer billings to date 1,100,000 1,555,000 3,130,000
Collections to date 1,000,000 1,750,000 2,950,000

Calculate the gross profit / (loss) that should be recognized for 2020, 2021, and 2022 using the percentage completion method. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

2020 2021 2022
Gross profit / (loss) $ $ $
Answer #1

Solution

2020 2021 2022
Gross profit $                551,250 $           204,000 $                  (5,250)

Working

Working 2020 2021 2022
A Contract Price $ 3,130,000 $ 3,130,000 $ 3,130,000
B Cost Incurred to Date $ 544,250 $ 903,650 $ 2,380,000
C Estimated cost yet to be incurred to complete the contract $ 1,010,750 $ 801,350 $ 0
D = B+C Total Cost $ 1,555,000 $ 1,705,000 $ 2,380,000
E = (D/B) x 100 % of Completion 35% 53% 100%
F = A x E Revenue to date $ 1,095,500 $ 1,658,900 $ 3,130,000
G Revenue of Previous year $ 0 $ 1,095,500 $ 1,658,900
H = F - G Net Revenue this year $ 1,095,500 $ 563,400 $ 1,471,100
I Cost to date $ 544,250 $ 903,650 $ 2,380,000
J Cost to date of previous year $ 0 $ 544,250 $ 903,650
K = I - J Net Cost for the year $ 544,250 $ 359,400 $ 1,476,350
L = H - K Gross Profits $ 551,250 $ 204,000 -$ 5,250